# xBound: Join Size Lower Bounds

### v1

@@ Abstract.
Cloud database vendors invest substantial resources into their query optimizers, and for good reason. Cardinality estimation, a cornerstone of the optimizer, is critical for the selection of efficient query plans, as well as downstream tasks such as resource allocation and query scheduling. Yet, as many practitioners and researchers have noted, it is also the optimizer’s Achilles heel. Prior studies on a number of industrial-strength databases show substantial cardinality estimation errors on all tested systems, with a far greater tendency to underestimate than to overestimate. Unfortunately, cardinality underestimation is more problematic than overestimation, as it misleads the optimizer to choose plans designed for small data, leading to underprovisioned CPU and memory.
While previous work on pessimistic cardinality estimation has proposed provable join size upper bounds, such methods can only correct overestimation, leaving the more harmful problem of underestimation unaddressed. To fill this critical gap, we introduce xBound, the very first framework for deriving provable join size lower bounds. xBound successfully reduces underestimation in real systems: On the JOBlight benchmark, it corrects 17.5% of subexpression underestimates in DuckDB and 8.7% in PostgreSQL, while on a Microsoft enterprise workload, it fixes 36.1% of Fabric Data Warehouse’s underestimates, demonstrating a significant step towards solving this long-standing problem.
@@ 1. Introduction
Figure 1. Overview of cardinality bound violations across all join subexpressions in JOBlight, when estimating with DuckDB 1.4, PostgreSQL 18, and Fabric Data Warehouse.
@@
After five decades of research and development, query optimization remains an “unsolved” problem. As pointed out by both practitioners and researchers, cardinality estimation, one of the crucial pillars of the query optimizer, is its Achilles heel (Lohman, 2014, 2017; Leis et al., 2015, 2025). By estimating the size of intermediate results, cardinality estimation is the primary input to the cost model for choosing the best execution plan. Additionally, it is essential for resource management, such as allocating memory and CPU, and determining the degree of parallelism in distributed databases. A seminal analysis by Leis et al. (Leis et al., 2015) empirically evaluated PostgreSQL (Stonebraker and Rowe, 1986), HyPer (Kemper and Neumann, 2011), and three commercial databases, revealing that all tested database systems routinely produce large cardinality estimation errors and that join size estimation is particularly fragile.
@@
This challenge has spurred a wealth of research on cardinality estimation, spanning sketch-based, sampling-based, to ML-based methods (see recent experimental surveys (Harmouch and Naumann, 2017; Han et al., 2021; Heinrich et al., 2025)). However, no matter how well a method has been shown to perform empirically, there is rarely any guarantee on the correctness of the estimation. In response, the theory community has recently proposed provable bounds for cardinality estimation, most notably the recent LpBound (Zhang et al., 2025). Yet, all these proposals provide only upper bounds. Rather than serving as direct replacements for the cardinality estimator, these upper bounds are best suited for correcting the original optimizer’s overestimates (by capping the estimates). Crucially, however, the complementary problem, safeguarding against underestimates, remains unaddressed.
Motivation. Interestingly, as shown in the same analysis (Leis et al., 2015), cardinality underestimation is far more prevalent than overestimation in practice, and this tendency increases with the number of joins. Moreover, although both misestimations are problematic, we argue that underestimation is generally more dangerous. It tricks the optimizer to choose plans best for small data, such as preferring nested-loops joins over hash- or merge-joins, broadcast joins over shuffle joins, causing catastrophic plan choices and order-of-magnitude performance slowdown. Regarding resource management, underestimation risks runtime resource exhaustion, leading to out-of-memory errors or excessive disk spilling. Conversely, overestimation misleads the optimizer to select conservative plans that are safe for big data, as well as overallocating runtime resources. Although this leads to degraded performance and wasted resources, the consequences are seldom catastrophic.
In this paper, we address the critical problem of safeguarding cardinality underestimation, by introducing xBound, the very first theoretical framework for computing provable join size lower bounds.
@@
Technical Overview. xBound, exploits, similar to LpBound (and sketch-based estimators (Wang et al., 2023; Heddes et al., 2024)), the key fact that the size of a join is the dot-product of the join-key degree vectors $\mathbf{a}$ and $\mathbf{b}$ of two relations $A$ and $B$, respectively. Our approach relies on the following observation: Lower bounds require inequalities that bound from below this very dot product; we refer to these as reverse inequalities. Fortunately, reverse inequalities only require a handful of statistics: $\ell_{\infty}$—the maximum key frequency, $\ell_{2}$—the Euclidean norm of the degree vector, and $\ell_{-\infty}$–the minimum key frequency. On the other hand, they require positive input vectors. We will thus first lower bound the number of non-zero entries in the entry-wise product $\mathbf{a}\odot\mathbf{b}$, enabling their application. This step requires the presence of min/max-values on the corresponding join columns, which are readily available in database catalogs and modern file formats.
To get a feeling how insightful lower and upper bounds can be in a real system, consider Fig. 1. We ran xBound and LpBound on all join subexpressions of JOBlight and compared their results against the estimates produced by DuckDB 1.4, PostgreSQL 18, and Fabric Data Warehouse (DW). Notably, xBound can correct 14.9% of DuckDB’s estimates and 6.9% of PostgreSQL’s. In comparison, LpBound corrects 8.3% and 0.4% of the estimates in the two systems, respectively. Evidently, this dual-bound approach enables us to both detect and mitigate more misestimates during query optimization.
@@
  • 1.
    We introduce xBound, the first framework for provable, non-trivial join size lower bounds, filling a critical gap in the cardinality estimation literature. In its current form, xBound estimates multi-way joins over the same join key.
  • 2.
    We extend our theoretical framework to obtain lower bounds over filtered table scans, supporting equality and range predicates as well as conjunctions and disjunctions.
  • 3.
    We empirically evaluate the impact of xBound in detecting and correcting cardinality underestimates in DuckDB v1.4, PostgreSQL 18, and Fabric DW, across the JOBlight (Kipf et al., 2019b), JOBjoin (Zhang et al., 2025), and STATS-CEB (Han et al., 2021) benchmarks, and on a subset of a Microsoft enterprise workload, MSSalesJoin (Alotaibi et al., 2024).
Structure. The rest of the paper is structured as follows: We provide the preliminaries needed in §2, including $\ell_{p}$-norms and dot-product inequalities. Then, in §3, we detail xBound’s components. Subsequently, we extend our framework to support lower bounds on filtered table scans in §4. We continue with the evaluation of xBound’s lower bounds in §5, and then with a discussion in §6. Afterwards, we outline related work in §7 and conclude in §8.
@@ 2. Preliminaries
Moreover, query $Q$ is acyclic if its relations $R_{1},\ldots,R_{n}$ can be placed on the nodes of a tree such that, for each variable $X_{i}\in\texttt{Vars}(Q)$, the set of tree nodes containing $X_{i}$ forms a connected component. In addition, $Q$ is called Berge-acyclic if it is acyclic and any two relations share at most one variable. For instance, the three-way joins $J_{1}$ and $J_{2}$ are Berge-acyclic, yet $J_{3}$ is cyclic:
@@
In this work, we focus solely on full Berge-acyclic conjunctive queries, with the additional restriction that joins are over the same join key. Notably, this is the case for many subexpressions of the well-known benchmarks, such as JOB (Leis et al., 2015; Kipf et al., 2019b) and STATS-CEB (Han et al., 2021).
@@ 2.1. Rearrangement Inequality
Before we move on to the dot-product reverse inequalities that we will be using, let us consider an inequality that connects dot-products of arbitrarily ordered vectors with dot-products of ordered sequences. This will turn useful for the overall understanding of xBound (and its difference to LpBound).
@@
In a nutshell, the RHS of the inequality is an upper bound on the dot product, while the LHS is a lower bound. In fact, the RHS is used in SafeBound (Deeds et al., 2023b, a), a pre-runner of LpBound (Zhang et al., 2025) (see  §7 for an exposition). What we (implicitly) leverage in xBound is the LHS, i.e., taking the dot-product between reversely ordered sequences. Yet, similar to LpBound, we will be directly bounding the dot-product by $\ell_{p}$-norms only.
2.2. Dot-Product Reverse Inequalities
As motivated in the introduction, we need the so-called reverse inequalities on the dot product of two degree vectors $\mathbf{a}$ and $\mathbf{b}$. The caveat is that this class of inequalities requires the input vectors to be strictly positive, i.e., $a_{i},b_{i}>0,\forall i\in[d]$. Recall that degree vectors are generally only non-negative, i.e., they contain zeros for non-existing keys in a relation. In  §3.1, we show how we can still apply these inequalities by lower-bounding the number of non-zero entries in both vectors.
We start with a well-known reverse inequality:
Inequality 2 (Pólya–Szegő Inequality (Pólya and Szegő, 1925)).
@@ 2.2. Dot-Product Reverse Inequalities
Let us inspect this inequality. To lower-bound the dot product $\mathbf{a}\cdot\mathbf{b}$, it requires the following vector norm statistics:
  • $\ell_{2}$: $\|a\|_{2}$ and $\|b\|_{2}$,
@@
As we will estimate the cardinalities of $k$-way joins, we also need the generalization of the above inequality for $k$ vectors. Fortunately, there is one. This kind of inequalities forms the interesting class of reverse Hölder’s inequalities. As a backbone, we will be using the (currently) tightest inequality in the literature, which indeed generalizes Ineq. (2). Hence, it infers the limitation of relying on $\ell_{2}$ only. We discuss possible extensions in  §6.
@@
Inequality 3 (Generalized Reversed Hölder’s Inequality (Krasopoulos and Bougoffa, 2022)).
@@
First, see that for $n=2$, the above yields the Pólya–Szegő inequality, Ineq. (2). Second, note that Ineq. (3) is permutation-sensitive, i.e., it does matter in which order the vectors are input. To see why this is the case, consider the $B_{k}$ constants: they consider only the statistics of the first $k$ vectors.
@@
To obtain a better lower bound in Ineq. (3), we will consider all (non-redundant) permutations of the input vectors, namely the (n1)!(n-1)! permutations obtained after excluding those obtained by swapping the first two elements, and take the maximum of the resulting lower bounds. While this increases the optimization time, this leads to better bounds. We briefly mention that it is even possible to have a convolution-like bound, by resorting to a subset dynamic programming in time $O(3^{n})$; however, we have not observed any improvement in the bounds.
Figure 3. xBound at runtime: Given a multi-way join on the same key, it (a) lower-bounds the number of joining keys (= $m$), (b) uses $m$ as the prefix length for evaluating degree-sequence norms on the base tables, and (c) takes the best bound from the generalized reverse Hölder’s inequality, min-degree bound, and $m$. The output represents a hard lower bound on the query size.
@@ 3. Join Size Lower Bounds
Overview. We can visualize xBound at runtime as in Fig. 3: First, (a) we infer a lower bound on the number of distinct joining keys. This is essentially the length of the degree sequences which we should take the norms from. Then, (b) we estimate the norms needed for the reverse inequalities in  §2.2, such as $\ell_{1}$, $\ell_{2}$, and $\ell_{\infty}$. By estimating, we mean computing a strict lower and upper bound of them. Finally, (c) we lower-bound the dot product by taking the maximum of all bounds output by the set of inequalities. Note that the value from (a) is also a valid lower bound.
Next, we start with the very first step: lower-bounding the number of joining keys.
3.1. Lower-Bounding $\ell_{0}$
@@
Our key insight is that we can first infer a lower bound on the number of joining keys, essentially $\ell_{A\,\bowtie\,B,\,X,\,0}^{-}$, and then apply the inequalities with norms confined to the first $\ell_{A\,\bowtie\,B,\,X,\,0}^{-}$ entries in the degree sequences of both relations. Note that this alone is a valid lower bound on the join size. As we will later show, dot-product inequalities can boost it even further.
Two-Way Joins. Let $\mathcal{K}_{A}$ and $\mathcal{K}_{B}$ be the set of keys in $A(X)$ and $B(X)$, respectively. Bounding the number of joining keys is equivalent to bounding $|\mathcal{K}_{A}\cap\mathcal{K}_{B}|$. Surely, having at our disposal only the cardinalities of the two sets, $\mathcal{K}_{A}$ and $\mathcal{K}_{A}$, we cannot have a better lower bound than 0. The key realization is that we can use the min/max-values of the columns, commonly referred to as zonemaps (Graefe, 2009), which are a standard feature in modern file formats (3; M. Kuschewski, D. Sauerwein, A. Alhomssi, and V. Leis (2023); 37; A. Afroozeh and P. Boncz (2025)). Let us understand why this changes the whole picture. Note that we can re-write $|\mathcal{K}_{A}\cap\mathcal{K}_{B}|$ as follows:
@@
Note that real-world systems typically do not store the exact values of $|\mathcal{K}_{A}|$ and $|\mathcal{K}_{B}|$. For example, an optimizer may rely on sketches such as HyperLogLog (Flajolet et al., 2007) to estimate the number of distinct values in a column. In its current form, xBound computes these numbers exactly, separately from the system. We get back to this discussion later, in  §6.
@@
To this end, we will be employing a lightweight generalization of the two-way join case. Denote the universe as $U\vcentcolon=\bigcup_{i}^{n}\mathcal{K}_{i}$. Then, it holds:
@@
(6) $=\sum_{i=1}^{n}|\mathcal{K}_{i}|-(n-1)|U|.$
Note that, for the case $n=2$, we obtain exactly the two-way join case, Eq. (4).
This concludes our approach to lower bounding the number of joining keys for an arbitrary number of joins on the same join key. In a subsequent section,  §4.2, we will further develop it, by partitioning the value range into equi-width bins, and applying the above strategy within each bin. Until then, we will first address the aspect of how to leverage this lower bound to enable reverse dot-product inequalities, which will further boost this value.

### v2

# The Case for Cardinality Lower Bounds

@@ Abstract.
Despite decades of research, cardinality estimation remains the optimizer’s Achilles heel, with industrial-strength systems exhibiting a systemic tendency toward underestimation. At cloud scale, this is a severe production vulnerability: in Microsoft’s Fabric Data Warehouse (DW), a mere 0.05% of extreme underestimates account for 95% of all CPU under-allocation, causing preventable slowdowns for thousands of queries daily. Yet recent theoretical work on provable upper bounds only corrects overestimation, leaving the more harmful problem of underestimation unaddressed. We argue that closing this gap is an urgent priority for the database community.
As a vital step toward this goal, we introduce xBound, the first theoretical framework for computing provable join size lower bounds. By clipping the optimizer’s estimates from below, xBound offers strict mathematical safety nets demanded by production systems—using only a handful of lightweight base table statistics. We demonstrate xBound’s practical impact on Fabric DW: on the StackOverflow-CEB benchmark, it corrects 23.6% of Fabric DW’s underestimates, yielding end-to-end query speedups of up to 20.1x, demonstrating that even a first step toward provable lower bounds can deliver meaningful production gains and motivating the community to further pursue this critical, open direction.
@@ 1. Introduction
Figure 1. Cardinality underestimation on the JOBlight and StackOverflow-CEB benchmarks. We correct the underestimates by clipping the optimizer’s estimates with xBound’s cardinality lower bounds.
@@
After five decades of research and development, query optimization remains an “unsolved” problem. As pointed out by both practitioners and researchers, cardinality estimation is the Achilles heel (Lohman, 2014, 2017; Leis et al., 2015, 2025) of query optimization. By estimating the size of intermediate results, cardinality estimation is the primary input to the cost model for plan selection. It is equally essential for resource management, dictating memory and CPU allocation. A seminal analysis by Leis et al. (Leis et al., 2015) empirically evaluated five open-source and commercial databases, revealing that all routinely produce large estimation errors, with join size estimation being particularly fragile.
@@
This challenge has spurred a wealth of research on cardinality estimation, spanning sketch-based, sampling-based, to ML-based methods (Harmouch and Naumann, 2017; Han et al., 2021; Heinrich et al., 2025). However, no matter how well a method performs empirically, there is rarely any guarantee on the correctness of the estimation. In response, the theory community has recently proposed provable bounds for cardinality estimation, most notably the recent LpBound (Zhang et al., 2025). Yet, all these proposals provide only upper bounds. Rather than serving as direct replacements for the cardinality estimator, these upper bounds are best suited for correcting the original optimizer’s overestimates (by capping the estimates). Crucially, however, the complementary problem, safeguarding against underestimates, remains unaddressed.
The Still Open Problem of Underestimation. The prior focus on upper bounds is fundamentally misaligned with production reality: cardinality underestimation is far more prevalent than overestimation in practice (Leis et al., 2015). When testing DuckDB v1.4, PostgreSQL18, and Fabric Data Warehouse (DW) on JOBlight (Kipf et al., 2019b) and StackOverflow-CEB (Han et al., 2021) [^1] That is, STATS-CEB run on the 220 GB StackOverflow dump from SQLStorm (Schmidt et al., 2025)., all three systems overwhelmingly underestimate result sizes (see Fig. 1). Moreover, although misestimation in both directions is problematic, underestimation is generally more dangerous. It tricks the optimizer into selecting fragile plans designed for small data, favoring nested-loops over hash joins or broadcast over shuffle joins, causing catastrophic plan regressions and order-of-magnitude slowdowns. At runtime, it risks resource starvation: out-of-memory errors and excessive disk spilling. Overestimation, by contrast, leads to overly conservative plans and overallocated resources, inefficient but seldom catastrophic.
The Peril of Extreme Underestimation. In large-scale production environments, extreme underestimation is a systemic vulnerability that severely degrades query performance. This is clearly signaled by CPU resource under-allocation. Fabric DW, for instance, allocates CPU per query stage based on its operators’ estimated output cardinality and row size. Fig. 2 reveals a heavy negative tail in Fabric’s CPU estimation: while ¿99% of stages face no underestimation (cardinality underestimation does not always lead to CPU underestimation), 1 in 10’000 stages suffers from ¿70x CPU under-allocation, with extreme outliers starved by over a thousandfold. At Fabric’s scale, this causes thousands of queries to suffer severe, preventable slowdowns every day. The database community can no longer afford to ignore this blind spot; it is imperative that both academia and industry actively shift their focus toward mitigating extreme underestimation.
The Case for Cardinality Lower Bounds. As a vital step towards solving this problem, we introduce xBound, the first theoretical framework for computing provable join size lower bounds using only a handful of lightweight base table statistics. Clipping optimizer estimates with these bounds significantly improves Q-errors, as shown in Fig. 1, notably reducing Fabric DW’s 90th-percentile Q-error of its underestimates on StackOverflow-CEB by 35.8x. Theoretical lower bounds are uniquely suited for adoption in production systems because they provide strict mathematical guarantees against catastrophic worst-case scenarios, unlike empirical methods that may perform well on average but lack necessary safety nets.
@@
Similar to LpBound and sketch-based estimators (Wang et al., 2023; Heddes et al., 2024), xBound the key fact that the size of a join is the inner product of the join-key degree vectors $\mathbf{a}$ and $\mathbf{b}$ of two relations $A$ and $B$, respectively. Our approach relies on the following observation: Lower bounds require inequalities that bound from below this very inner product; we refer to these as reverse inequalities. Fortunately, reverse inequalities only require a handful of statistics: $\ell_{\infty}$—the maximum key frequency, $\ell_{2}$—the Euclidean norm of the degree vector, and $\ell_{-\infty}$—the minimum key frequency. On the other hand, they require positive input vectors. We will thus first lower bound the number of non-zero entries in the entry-wise product $\mathbf{a}\odot\mathbf{b}$, enabling their application.
Results show that xBound corrects 23.6% of underestimates on the StackOverflow-CEB benchmark, yielding end-to-end query speedups of up to 20.1x on Fabric DW.
@@
  • (1)
    We make the case for cardinality lower bounds by pointing out severe underestimation in production, and introduce xBound, the first framework for provable lower bounds on join sizes. In its current form, xBound estimates multi-way joins over the same join key.
  • (2)
    We extend our framework to obtain lower bounds over filtered table scans, supporting equality and range predicates as well as conjunctions and disjunctions.
  • (3)
    We empirically evaluate the impact of correcting Fabric DW’s underestimates with xBound’s lower bounds on the StackOverflow-CEB benchmark.
Structure. The rest of the paper is structured as follows: We provide the preliminaries needed in §2, including $\ell_{p}$-norms and inner-product inequalities. Then, in §3, we detail xBound’s components. Subsequently, we extend our framework to support lower bounds on filtered table scans in §4. We continue with the evaluation of xBound’s lower bounds in §5, and then with a discussion in §6, where we motivate further research on lower bounds. Afterwards, we outline related work in §7 and conclude in §8.
Figure 2. CPU deficit due to cardinality underestimation of query stages executed in production of Fabric Data Warehouse in Feb. 2026 across all cloud regions.
@@ 2. Preliminaries
Query $Q$ is acyclic if its relations $R_{1},\ldots,R_{n}$ can be placed on the nodes of a tree such that, for each variable $X_{i}\in\texttt{Vars}(Q)$, the set of tree nodes containing $X_{i}$ forms a connected component. For instance, the three-way joins $J_{1}$ and $J_{2}$ are acyclic, [^3] Even Berge-acyclic, since every two relations share at most one variable. yet $J_{3}$ is cyclic:
@@
In this work, we focus solely on inner and outer joins on single keys. Notably, this is the case for many subexpressions of the well-studied benchmarks, such as JOB (Leis et al., 2015; Kipf et al., 2019b) and STATS-CEB (Han et al., 2021).
Join Pool. We refer to a join pool as a set of attributes that are joined over in the same query, e.g., the userid across all tables of a social network schema. Join pools can be discovered either directly from the schema or, in its absence, dynamically as the query workload is running. This definition will prove helpful when introducing the heavy partition in §3.5.
@@ 2.1. Rearrangement Inequality
Before we move on to the inner-product reverse inequalities that we will be using, let us consider an inequality that connects inner products of arbitrarily ordered vectors with inner products of ordered sequences. This will prove useful for the overall understanding of xBound, and its difference from LpBound (Zhang et al., 2025), a state-of-the-art framework for cardinality upper bounds.
@@
In a nutshell, the RHS of the inequality is an upper bound on the inner product, while the LHS is a lower bound. In fact, the RHS is used in SafeBound (Deeds et al., 2023b, a), a pre-runner of LpBound (Zhang et al., 2025) (see  §7 for an exposition). What we (implicitly) leverage in xBound is the LHS, i.e., taking the inner product between reversely ordered sequences. Yet, similar to LpBound, we will be directly bounding the inner product by $\ell_{p}$-norms only.
2.2. Inner-Product Reverse Inequalities
As motivated in the introduction, we need the so-called reverse inequalities on the inner product of two degree vectors $\mathbf{a}$ and $\mathbf{b}$. The caveat is that this class of inequalities requires the input vectors to be strictly positive, i.e., $a_{i},b_{i}>0,\forall i\in[d]$. Recall that degree vectors are generally only non-negative, i.e., they contain zeros for non-existing keys in a relation. In  §3.1, we show how we can still apply these inequalities by lower-bounding the number of non-zero entries in both vectors. We start with an intuitive reverse inequality:
Inequality 2 (Min-Degree Inequality).
$0<m_{a}\leq a_{i}<\infty,\;0<m_{b}\leq b_{i}<\infty,\;\forall i\in[d].$
Then
$\sum_{i=1}^{d}a_{i}b_{i}\geq\max(m_{a}\cdot\displaystyle\sum_{i=1}^{d}b_{i},m_{b}\cdot\displaystyle\sum_{i=1}^{d}a_{i}).$
This inequality tells us that, to lower-bound the inner product, it suffices to assume that one vector is filled with its smallest element. Notably, this requires only the vector’s element sum and its smallest element. Next, we continue with a more involved reverse inequality:
Inequality 3 (Pólya–Szegő Inequality (Pólya and Szegő, 1925)).
@@ 2.2. Inner-Product Reverse Inequalities
Let us inspect this inequality. To lower-bound the inner product $\mathbf{a}\cdot\mathbf{b}$, it requires the following vector norm statistics:
  • $\ell_{2}$: $\|\mathbf{a}\|_{2}$ and $\|\mathbf{b}\|_{2}$,
@@
As we will estimate the cardinalities of $k$-way joins, we also need the generalization of the above inequality for $k$ vectors. Fortunately, there is one. Such inequalities form the interesting class of reverse Hölder’s inequalities. As a backbone, we will be using the (currently) tightest inequality in the literature, which indeed generalizes Ineq. (3). Hence, it infers the limitation of relying on $\ell_{2}$ only. We discuss possible extensions in  §6.
@@
Inequality 4 (Generalized Reverse Hölder’s Inequality (Krasopoulos and Bougoffa, 2022)).
@@
First, see that for $n=2$, the above reduces to the Pólya–Szegő inequality, Ineq. (3). Second, note that Ineq. (4) is permutation-sensitive, i.e., it does matter in which order the vectors are input. To see why this is the case, consider the $B_{k}$ constants: they consider only the statistics of the first $k$ vectors.
@@
To obtain a better lower bound in Ineq. (4), we will consider all (non-redundant) permutations of the input vectors, namely the (n1)!(n-1)! permutations obtained after excluding those that differ only by swapping the first two elements, and take the maximum of the resulting lower bounds. While this increases the optimization time, this leads to better bounds. We briefly mention that it is even possible to have a convolution-like bound, by resorting to a subset dynamic program in time $O(3^{n})$; however, we have not observed any improvement in the bounds.
Figure 4. xBound’s estimation method: Given a multi-way join on the same key, it (a) lower-bounds the number of joining keys (= $m$), (b) uses $m$ as the prefix length for evaluating degree-sequence norms on the base tables, and (c) takes the best bound from the generalized reverse Hölder’s inequality, min-degree bound, and $m$. The output represents a lower bound on the query size. Note that this visualization does not cover the heavy partition (§3.5).
@@ 3. Join Size Lower Bounds
Overview. We visualize xBound’s estimation strategy in Fig. 4: First, (a) we infer a lower bound on the number of distinct joining keys. This is essentially the length of the degree sequences which we should take the norms from. Then, (b) we estimate the norms needed for the reverse inequalities in  §2.2, such as $\ell_{1}$, $\ell_{2}$, and $\ell_{\infty}$. By estimating, we mean computing a strict lower and upper bound of them. Finally, (c) we lower-bound the inner product by taking the maximum of all bounds output by the set of inequalities. Note that the value from (a) is also a valid lower bound.
Next, we start with the very first step: obtaining hard lower bounds the number of distinct joining keys.
3.1. Hard $\ell_{0}$ Lower Bounds
@@
Our key insight is that we can first infer a lower bound on the number of joining keys, essentially $\ell_{A\,\bowtie\,B,\,X,\,0}^{-}$, and then apply the inequalities with norms confined to the first $\ell_{A\,\bowtie\,B,\,X,\,0}^{-}$ entries in the degree sequences of both relations. Note that this alone is a valid lower bound on the join size. As we will later show, inner-product inequalities can boost it even further.
Two-Way Joins. Let $\mathcal{K}_{A}$ and $\mathcal{K}_{B}$ be the set of keys in $A(X)$ and $B(X)$, respectively. Bounding the number of joining keys is equivalent to bounding $|\mathcal{K}_{A}\cap\mathcal{K}_{B}|$. Surely, having at our disposal only the cardinalities of the two sets, $\mathcal{K}_{A}$ and $\mathcal{K}_{A}$, we cannot have a better lower bound than 0. The key realization is that we can use the min/max-values of the columns, commonly referred to as zonemaps (Graefe, 2009), which are a standard feature in modern file formats (par, 2025; Kuschewski et al., 2023; vor, 2025; Afroozeh and Boncz, 2025). Let us understand why this changes the whole picture. Note that we can re-write $|\mathcal{K}_{A}\cap\mathcal{K}_{B}|$ as follows:
@@
To this end, we will be employing a lightweight generalization of the two-way join case. Denote the universe as $U\vcentcolon=\bigcup_{i}^{n}\mathcal{K}_{i}$. Then:
@@
(6) $=\sum_{i=1}^{n}|\mathcal{K}_{i}|-(n-1)|U|,$
where we used the union bound in the second step. Note that, for the case $n=2$, we obtain exactly the two-way join case, Eq. (4).
3.2. Probabilistic $\ell_{0}$ Lower Bounds
The above section discussed how to get hard lower bounds on the number of distinct joining keys. However, we noticed that collecting the exact $\ell_{0}$ statistics can be slow, especially for larger tables. Several database systems, including Fabric DW, already store on their catalogs HyperLogLog (Flajolet et al., 2007) sketches of their columns. Indeed, these can be used to get a probabilistic lower bound on $\ell_{0}$, by plugging them in Eq. (6) for the $|\mathcal{K}_{i}|$ values. Alternatively, one can use the inclusion-exclusion principle, yet this leads to high-variant estimates (Dasgupta et al., 2016, Sec. 2.2)
In fact, there are specialized sketches that can estimate the intersection size of $n$ sets directly (Dasgupta et al., 2016; Ertl, 2021). To this end, we used the ThetaSketch (Dasgupta et al., 2016), as it is readily provided in Apache’s DataSketches library. Hence, instead of lower bounding $|\cap_{i=1}^{n}\mathcal{K}_{i}|$ via the union bound—recall Eq. (6), we can get a better lower bound, with a 99% confidence, by intersecting the Theta sketches corresponding to the $n$ sets. These sketches are stored as pre-computed base table statistics; we come later in §4 onto how we manage them.
Notably, since the inner-product inequalities are deterministic functions of the estimated intersection lower bound, the 99% confidence guarantee provided by the Theta sketch propagates unchanged to the derived query size lower bounds.
Outlook. This concludes our approach to lower-bounding the number of distinct joining keys for an arbitrary number of joins on the same join key. In a subsequent section (§4.2), we will further develop it, by partitioning the value range into equi-width bins, and applying the above strategy within each bin. Until then, we address the aspect of how to leverage this very lower bound to enable reverse inner-product inequalities, which will further boost it.